# Lead Times in Flux: Analyzing Airbnb Dynamics During Global Upheavals (2018-2022)

### v1

@@ 5 Discussion
Our analysis showed differential recovery patterns among the cities studied, contrasting the influence of local policies and destination-specific factors on tourism resilience and booking behaviors. Cities like Miami and Austin exhibited quicker recoveries in domestic travel lead times, returning closer to pre-pandemic patterns by 2022. This resilience may be attributed to less stringent local COVID-19 restrictions, a higher proportion of leisure travelers, and their appeal as destinations—factors that align with prior research indicating that destinations with flexible policies and strong domestic appeal recovered more swiftly from pandemic-induced downturns (Sigala, 2020). In contrast, Boston and San Francisco showed prolonged divergences in both domestic and international travel lead times, likely due to stricter local regulations, reliance on international business travel, and cautious reopening strategies. These observations are consistent with studies noting that cities dependent on international and business travel faced greater challenges during the pandemic (Götz et al., 2021).
These findings align with theories of risk perception and consumer behavior during crises, suggesting that uncertainty and perceived risk lead to more cautious consumer behaviors, such as delayed booking and a preference for flexible options (Lo et al., 2011). The increased lead times and heightened variability during the pandemic years corroborate this, particularly in international travel, which experienced greater disruptions due to heightened risk perceptions and fluctuating international policies. The impact of global crises on travel patterns has been extensively researched, and our study extends this understanding by showing how these factors influence not only overall visitation numbers but also the temporal dynamics of booking behaviors and forecasting accuracy.
Furthermore, the limitations of traditional summary statistics in capturing these behavioral shifts underscore the need for advanced analytical approaches in tourism research. By introducing the normalized L1 distance metric, our study contributes to methodological advancements in understanding complex consumer behaviors. This metric effectively quantified changes in booking lead times and showed shifts not apparent through traditional measures, enhancing our ability to understand and forecast traveler behavior in response to external shocks. As our toy example demonstrated, even small divergences captured by the normalized L1 distance can lead to significant under- or overestimation of demand. The derived forecast error bound offers a practical tool to assess and mitigate these errors by estimating the normalized L1 distance and adjusting forecasting models accordingly.
The sustained divergences in booking behaviors have important implications for the tourism and hospitality industry. Businesses should anticipate continued changes in lead time distributions and not expect a return to pre-pandemic patterns. These shifts suggest that traditional forecasting models based on historical data may no longer suffice, necessitating updates to demand forecasting and revenue management strategies. Practitioners relying on naive forecasting methods, such as the pickup method, should be aware of potential forecast errors due to shifts in lead time distributions. Moreover, the increasing trend in year-over-year divergences for international origin travel implies that booking behaviors are still evolving, underscoring the importance of flexibility and adaptability in business operations. Companies should monitor these trends closely to respond promptly to changes in traveler behavior, such as offering flexible booking policies or tailoring marketing efforts to match current planning horizons.
While our study offers valuable insights, it has limitations. The analysis is confined to four U.S. cities and one accommodation platform, which may limit generalizability. Future research could expand to include a broader range of destinations and accommodation types. Additionally, ongoing developments, such as changes in remote work policies and evolving traveler preferences, may continue to impact booking behaviors. Longitudinal studies extending beyond 2022 would help in understanding the long-term effects of the pandemic and other global events on tourism. Further methodological advancements could enhance the applicability of the normalized L1 distance metric and the derived forecast error bounds. Integrating the normalized L1 distance into predictive models or combining it with other metrics may offer more robust tools for industry stakeholders.
Acknowledgement
[…]

### v3

@@ 5 Discussion
5.1 Key Insights on Lead-Time Distributions
Our analysis shows that the COVID-19 pandemic triggered marked changes in Airbnb booking lead times, often in ways that simple mean or median metrics would fail to capture. By studying the entire distribution of lead times through the normalized L1 distance, we revealed substantial shifts, including cases where domestic travel in certain cities (e.g., Austin, Miami) rebounded more quickly, whereas others (e.g., Boston, San Francisco) continued to diverge from 2018 baselines. While these differences could reflect variations in local economic composition, traveler preferences, or policy approaches (Sigala, 2020a), our data do not directly measure the causal impact of local regulations. Rather, the persistent L1 divergences remind us that forecasting models relying heavily on historical norms may underperform when external shocks alter traveler behavior in fundamental ways.
5.1.1 Distribution-Wide Reallocation of Lead Times
Several city-year examples illustrate how L1 can capture internal shifts that remain obscured by stable or modest changes in averages. For instance, in San Francisco’s domestic travel during 2020, the mean lead time rose by only about 3% (see Table S5), yet the mean $l^{1}_{{\text{YoY2018}},t}$ exceeded 0.16 (Table S9), implying that a notable fraction of bookings migrated to the extremes (very short or very long lead times). Likewise, destination Miami’s domestic bookings in 2020 showed a mean that was nearly unchanged from 2018, yet the L1 distance spiked to around 0.11, indicating tail-driven reallocation. And in origin Boston’s domestic market in 2020, the median increased 3% from the previous year while the median $l^{1}_{{\text{YoY}},t}$ was .11 suggesting that while many travelers booked further out, a subset planned well in advance or much earlier than they did in the previous year, preserving considerable spread overall. In each case, L1 reveals the depth of these distributional shifts, underscoring how stable or even partially “recovered” means and medians can mask significant reconfigurations of booking horizons.
5.2 Implications for Forecasting and Revenue Management
In practical terms, elevated L1 distances highlight situations where the historical “booking curve” may be a poor guide to current or future patterns. Revenue managers adopting naive pickup methods could systematically under- or over-predict demand if more guests begin booking at very short notice, or conversely, further in advance. Mitigation strategies include:
  • Recalibrated Forecasts: Incorporate distribution-aware metrics into forecast models, periodically re-estimating how booking timelines have shifted.
  • Scenario Analyses: Use the L1 distance as a red flag to explore best- and worst-case demand scenarios, particularly when pivoting to new pricing or marketing campaigns.
  • Monitoring International vs. Domestic Patterns: As our results illustrate, cross-border travel can diverge more drastically from historical patterns, warranting special attention to the L1 divergences in that segment.
5.3 Forward-Looking Extensions of the L1 Distance
Although the present study focuses on final lead-time distributions, a natural extension is to track how in-progress (partial) bookings compare to the historical shape of bookings at the same horizon. Concretely:
  1. 1.
    Partial Distribution Construction: For each arrival date, identify what fraction of total bookings (so far) accrued at lead times from 00 to $H$ days out, where $H$ is a “checkpoint” (e.g., 30 days before arrival).
  2. 2.
    Comparative Baseline: Compute an average partial distribution at $H$ days out using data from prior years or seasons. This yields a historical reference $L^{(H)}_{\text{hist}}$ to which the current partial distribution $L^{(H)}_{t}$ can be compared.
  3. 3.
    Partial L1 Metric:
    $D_{H}(t)\;=\;\frac{1}{2}\sum_{\delta=0}^{H}|L^{(H)}_{t}(\delta)-L^{(H)}_{\text{hist}}(\delta)|.$
    A higher $D_{H}(t)$ signals that current booking behavior is unfolding quite differently from the norm.
  4. 4.
    Early Warning: When $D_{H}(t)$ exceeds a chosen threshold, managers can anticipate a final distribution that may deviate from historical patterns. They might respond by adjusting prices, experimenting with targeted promotions, or revising inventory allocations.
Because these “forward-looking” metrics capture divergences before the final booking curve is observed, they could serve as an early-warning tool in times of heightened volatility. While we have not implemented this approach in the current analysis, it represents a natural follow-up for practitioners seeking near-real-time anomaly detection.
5.4 Conclusion and Future Directions
By emphasizing the full distribution of lead times, this study demonstrates how the COVID-19 pandemic produced lasting changes that are not fully captured by traditional statistics such as the mean or median. In revealing reallocation among short-, medium-, and long-lead segments, the normalized L1 distance emerges as a valuable diagnostic tool for detecting and quantifying booking-behavior shifts. These insights are particularly relevant to industry stakeholders looking to adapt forecasting and operational strategies to volatile travel markets.
Despite this contribution, our analysis is subject to several limitations. First, we focused on four major U.S. cities using data from a single accommodation platform. Results may not generalize to all markets or lodging segments, particularly those with smaller or more heterogeneous demand bases. Second, we did not isolate other potential drivers of lead-time divergence, such as flight capacity, local pandemic severity, or remote-work policies. Incorporating these factors might clarify whether certain market segments or traveler demographics are more prone to last-minute shifts.
Building on these insights, future research could broaden the dataset to include additional destinations or compare multiple lodging types. Methodologically, integrating the L1 distance into machine learning or advanced time-series models could improve forecast robustness, especially under fluctuating conditions. Another promising direction involves piloting a forward-looking or partial-L1 approach in real-world revenue management systems, where even small early warnings about distributional shifts can yield significant benefits in pricing, marketing, and inventory controls.
Data Availability
The primary dataset used in our study, Lead Times in Flux: Analyzing Airbnb Dynamics During Global Upheavals (2018-2022), is not publicly available due to confidentiality constraints.
Supplementary Materials
5.5 Derivation of the Forecast Error Bound
We aim to derive the bound:
$|\epsilon|\leq\frac{2D\left(1-\frac{\Delta}{\Delta_{\text{max}}}\right)}{C_{\text{hist}}(\Delta)}$
Definitions
Let:
  • $L_{\text{hist}}(k)$ and $L_{\text{actual}}(k)$ be the historical and actual lead time distributions.
  • $C_{\text{hist}}(\Delta)=\sum_{k=1}^{\Delta}L_{\text{hist}}(k)$ and $C_{\text{actual}}(\Delta)=\sum_{k=1}^{\Delta}L_{\text{actual}}(k)$ be their cumulative distributions.
  • $D=\frac{1}{2}\sum_{k=1}^{\Delta_{\text{max}}}|L_{\text{hist}}(k)-L_{\text{actual}}(k)|$ be the normalized L1 distance.
  • $\Delta_{\text{max}}$ be the maximum lead time.
The relative forecast error $\epsilon$ is defined as:
$\epsilon=\frac{\hat{B}_{\text{total}}-B_{\text{total}}}{B_{\text{total}}}$
where:
  • $\hat{B}_{\text{total}}=\frac{B_{\text{obs}}}{C_{\text{hist}}(\Delta)}$ is the forecasted total bookings using the pickup method.
  • $B_{\text{obs}}=C_{\text{actual}}(\Delta)B_{\text{total}}$ is the actual number of bookings observed up to lead time $\Delta$.
Derivation
Substitute $\hat{B}_{\text{total}}$ and $B_{\text{obs}}$ into $\epsilon$:
$\epsilon=\frac{\frac{B_{\text{obs}}}{C_{\text{hist}}(\Delta)}-B_{\text{total}}}{B_{\text{total}}}=\frac{\frac{C_{\text{actual}}(\Delta)B_{\text{total}}}{C_{\text{hist}}(\Delta)}-B_{\text{total}}}{B_{\text{total}}}$
Simplify:
$\epsilon=\frac{C_{\text{actual}}(\Delta)}{C_{\text{hist}}(\Delta)}-1=\frac{C_{\text{actual}}(\Delta)-C_{\text{hist}}(\Delta)}{C_{\text{hist}}(\Delta)}$
Take the absolute value:
$|\epsilon|=\left|\frac{C_{\text{actual}}(\Delta)-C_{\text{hist}}(\Delta)}{C_{\text{hist}}(\Delta)}\right|$
We need to bound $|C_{\text{actual}}(\Delta)-C_{\text{hist}}(\Delta)|$. Consider:
$C_{\text{actual}}(\Delta)-C_{\text{hist}}(\Delta)=\sum_{k=1}^{\Delta}(L_{\text{actual}}(k)-L_{\text{hist}}(k))$
Since the total difference between the distributions is $2D$:
$\sum_{k=1}^{\Delta_{\text{max}}}|L_{\text{actual}}(k)-L_{\text{hist}}(k)|=2D$
The sum from $k=1$ to $\Delta$ accounts for a fraction $\frac{\Delta}{\Delta_{\text{max}}}$ of the total possible divergence (assuming the divergence is uniformly distributed over $\Delta_{\text{max}}$).
Therefore, the maximum cumulative difference up to $\Delta$ is:
$|C_{\text{actual}}(\Delta)-C_{\text{hist}}(\Delta)|\leq 2D\left(1-\frac{\Delta}{\Delta_{\text{max}}}\right)$
Substituting back into $|\epsilon|$:
$|\epsilon|\leq\frac{2D\left(1-\frac{\Delta}{\Delta_{\text{max}}}\right)}{C_{\text{hist}}(\Delta)}$
This completes the derivation of the forecast error bound.
Implications
  • As $\Delta$ approaches $\Delta_{\text{max}}$, $\left(1-\frac{\Delta}{\Delta_{\text{max}}}\right)$ approaches zero, causing the bound $|\epsilon|$ to approach zero.
  • As $D\rightarrow 1$, the bound becomes less informative.
Supplementary Figures and Tables
Figure S1: A heatmap displaying the correlations between L1 Yo2018 divergences for Austin, Boston, Miami, and San Francisco and travel types (domestic vs. international), with colder colors indicating stronger positive correlations in divergence behavior.
Figure S2: A heatmap displaying the correlations between L1 YoY divergences for Austin, Boston, Miami, and San Francisco and travel types (domestic vs. international), with colder colors indicating stronger positive correlations in divergence behavior.
Figure S3: Seasonal-Trend decomposition (STL) of the year-over-2018 ($l^{1}_{{\text{YoY2018}},t}$) divergence time series for destination lead times in four U.S. cities (Austin, Boston, Miami, San Francisco). Each panel corresponds to one city-route combination (e.g., Austin-domestic, Boston-international), and the $l^{1}_{{\text{YoY2018}},t}$ values are plotted monthly from 2019 to 2023 on the horizontal axis. The blue curve (Trend) shows the longer-term evolution of divergence from the 2018 baseline, the red curve (Seasonal) captures recurring seasonal fluctuations (e.g., holiday or summer travel effects), and the green curve (Remainder) reflects short-term or irregular variations not explained by trend or seasonality.
Figure S4: Seasonal-Trend decomposition (STL) of the year-over-2018 ($l^{1}_{{\text{YoY2018}},t}$) divergence time series for origin lead times in four U.S. cities (Austin, Boston, Miami, San Francisco). As in Figure S3, each panel shows monthly $l^{1}_{{\text{YoY2018}},t}$ from 2019 to 2023 for a specific city-route combination (e.g., Miami-international). The decomposition splits the overall divergence (Value, on the vertical axis) into: a blue Trend line indicating the long-run trajectory of how far lead-time distributions have shifted from their 2018 baseline; a red Seasonal line capturing repeating patterns driven by annual or event-related cycles; and a green Remainder line revealing short-lived or irregular changes.
Figure S5: Seasonal-Trend decomposition (STL) of the year-over-year ($l^{1}_{{\text{YoY}},t}$) divergence time series for destination lead times in four U.S. cities (Austin, Boston, Miami, San Francisco). Each panel corresponds to one city-route combination (e.g., Austin-domestic, Boston-international) and spans monthly data from 2019 through 2023. The blue Trend curve tracks the longer-term evolution in how destination lead-time distributions diverge from the previous year, capturing sustained increases or decreases over time. The red Seasonal curve isolates recurring, typically annual patterns (e.g., seasonal peaks tied to holidays or summer travel). The green Remainder curve represents short-lived or irregular fluctuations unaccounted for by the trend and seasonal components.
Figure S6: Seasonal-Trend decomposition (STL) of the year-over-year ($l^{1}_{{\text{YoY}},t}$) divergence time series for origin lead times in four U.S. cities (Austin, Boston, Miami, San Francisco). As in Figure S5, each panel represents a specific city-route pairing over monthly data from 2019 through 2023. The decomposition splits the observed YoY divergence into a blue Trend component, a red Seasonal component, and a green Remainder component. The Trend line reflects longer-term shifts in how far origin lead times deviate from the prior year’s norms, the Seasonal line highlights periodic patterns linked to factors like school breaks or major holidays, and the Remainder captures one-off spikes or dips not explained by either trend or seasonality.
city year corridor mean median sd
Austin 2018 destination 44 28 49
Austin 2019 destination 46 29 50
Austin 2020 destination 46 28 52
Austin 2021 destination 49 34 49
Austin 2022 destination 49 30 56
Austin 2018 origin 45 28 50
Austin 2019 origin 46 29 51
Austin 2020 origin 44 27 53
Austin 2021 origin 49 33 50
Austin 2022 origin 53 34 57
Boston 2018 destination 47 29 53
Boston 2019 destination 49 30 55
Boston 2020 destination 59 35 65
Boston 2021 destination 57 40 55
Boston 2022 destination 65 46 63
Boston 2018 origin 52 32 57
Boston 2019 origin 54 34 59
Boston 2020 origin 56 34 62
Boston 2021 origin 59 41 57
Boston 2022 origin 67 45 68
Miami 2018 destination 39 24 46
Miami 2019 destination 40 24 46
Miami 2020 destination 39 22 49
Miami 2021 destination 39 26 43
Miami 2022 destination 43 26 51
Miami 2018 origin 35 20 44
Miami 2019 origin 36 19 46
Miami 2020 origin 33 17 45
Miami 2021 origin 39 23 46
Miami 2022 origin 40 21 52
San Francisco 2018 destination 52 37 50
San Francisco 2019 destination 52 36 52
San Francisco 2020 destination 54 32 58
San Francisco 2021 destination 49 36 47
San Francisco 2022 destination 61 46 55
San Francisco 2018 origin 42 27 47
San Francisco 2019 origin 45 28 50
San Francisco 2020 origin 43 28 48
San Francisco 2021 origin 50 36 47
San Francisco 2022 origin 55 37 56
Table S1: Airbnb summary lead times domestic: Summary statistics for domestic Airbnb bookings from 2018 to 2022, including mean, median, and standard deviation of lead times across four U.S. cities (Austin, Boston, Miami, and San Francisco). Origin/destination indicates whether the booking was from/to the city.
city year corridor mean median sd
Austin 2018 destination 60 41 60
Austin 2019 destination 61 42 60
Austin 2020 destination 77 52 74
Austin 2021 destination 48 31 50
Austin 2022 destination 50 32 53
Austin 2018 origin 56 38 57
Austin 2019 origin 58 38 60
Austin 2020 origin 71 43 74
Austin 2021 origin 50 32 54
Austin 2022 origin 60 41 61
Boston 2018 destination 61 43 58
Boston 2019 destination 64 45 62
Boston 2020 destination 95 74 80
Boston 2021 destination 64 46 60
Boston 2022 destination 70 54 62
Boston 2018 origin 57 39 58
Boston 2019 origin 59 40 59
Boston 2020 origin 72 45 73
Boston 2021 origin 50 32 55
Boston 2022 origin 61 42 62
Miami 2018 destination 61 43 59
Miami 2019 destination 61 42 60
Miami 2020 destination 78 54 75
Miami 2021 destination 41 26 50
Miami 2022 destination 50 32 55
Miami 2018 origin 45 27 51
Miami 2019 origin 45 26 52
Miami 2020 origin 48 25 60
Miami 2021 origin 39 23 47
Miami 2022 origin 46 27 53
San Francisco 2018 destination 67 49 60
San Francisco 2019 destination 69 50 62
San Francisco 2020 destination 93 71 79
San Francisco 2021 destination 55 37 57
San Francisco 2022 destination 64 48 56
San Francisco 2018 origin 50 34 53
San Francisco 2019 origin 53 35 55
San Francisco 2020 origin 59 35 65
San Francisco 2021 origin 44 29 49
San Francisco 2022 origin 56 38 57
Table S2: Airbnb summary lead times international: Summary statistics for international Airbnb bookings from 2018 to 2022, including mean, median, and standard deviation of lead times across four U.S. cities (Austin, Boston, Miami, and San Francisco). Origin/destination indicates whether the booking was from/to the city.
city year corridor mean median sd
Austin 2019 destination 1.036 1.040 1.025
Austin 2020 destination 0.998 0.962 1.041
Austin 2021 destination 1.078 1.204 0.951
Austin 2022 destination 1.007 0.897 1.125
Austin 2019 origin 1.022 1.028 1.020
Austin 2020 origin 0.961 0.911 1.026
Austin 2021 origin 1.096 1.251 0.958
Austin 2022 origin 1.097 1.037 1.141
Boston 2019 destination 1.043 1.038 1.038
Boston 2020 destination 1.198 1.164 1.178
Boston 2021 destination 0.960 1.153 0.848
Boston 2022 destination 1.145 1.131 1.142
Boston 2019 origin 1.046 1.057 1.038
Boston 2020 origin 1.021 0.977 1.049
Boston 2021 origin 1.064 1.228 0.927
Boston 2022 origin 1.132 1.084 1.178
Miami 2019 destination 1.039 1.043 1.022
Miami 2020 destination 0.973 0.903 1.038
Miami 2021 destination 0.999 1.165 0.888
Miami 2022 destination 1.109 0.990 1.191
Miami 2019 origin 1.006 0.972 1.049
Miami 2020 origin 0.928 0.913 0.977
Miami 2021 origin 1.169 1.349 1.020
Miami 2022 origin 1.034 0.910 1.120
San Francisco 2019 destination 1.000 0.969 1.032
San Francisco 2020 destination 1.023 0.902 1.130
San Francisco 2021 destination 0.918 1.105 0.799
San Francisco 2022 destination 1.245 1.285 1.184
San Francisco 2019 origin 1.052 1.048 1.053
San Francisco 2020 origin 0.948 0.979 0.957
San Francisco 2021 origin 1.176 1.310 0.987
San Francisco 2022 origin 1.104 1.024 1.194
Table S3: Year-over-Year (YoY) metrics for domestic travel. The table presents the mean, median, and standard deviation (sd) of lead times for Airbnb bookings across all months for Austin, Boston, Miami, and San Francisco from 2019 to 2022.
city year corridor mean median sd
Austin 2019 destination 1.008 1.013 1.004
Austin 2020 destination 1.271 1.245 1.225
Austin 2021 destination 0.617 0.593 0.681
Austin 2022 destination 1.048 1.037 1.059
Austin 2019 origin 1.029 1.009 1.052
Austin 2020 origin 1.227 1.115 1.239
Austin 2021 origin 0.697 0.757 0.723
Austin 2022 origin 1.205 1.254 1.129
Boston 2019 destination 1.057 1.049 1.065
Boston 2020 destination 1.482 1.659 1.297
Boston 2021 destination 0.676 0.626 0.748
Boston 2022 destination 1.094 1.154 1.038
Boston 2019 origin 1.019 1.014 1.022
Boston 2020 origin 1.233 1.137 1.242
Boston 2021 origin 0.697 0.717 0.757
Boston 2022 origin 1.214 1.287 1.113
Miami 2019 destination 1.006 0.991 1.025
Miami 2020 destination 1.281 1.281 1.238
Miami 2021 destination 0.526 0.477 0.666
Miami 2022 destination 1.220 1.230 1.113
Miami 2019 origin 0.999 0.985 1.010
Miami 2020 origin 1.082 0.958 1.159
Miami 2021 origin 0.801 0.916 0.787
Miami 2022 origin 1.179 1.179 1.118
San Francisco 2019 destination 1.020 1.013 1.040
San Francisco 2020 destination 1.361 1.423 1.262
San Francisco 2021 destination 0.584 0.521 0.720
San Francisco 2022 destination 1.165 1.298 0.995
San Francisco 2019 origin 1.041 1.029 1.050
San Francisco 2020 origin 1.123 1.006 1.173
San Francisco 2021 origin 0.751 0.836 0.757
San Francisco 2022 origin 1.255 1.306 1.158
Table S4: Year-over-Year (YoY) metrics for international travel. The table presents the mean, median, and standard deviation (sd) of lead times for Airbnb bookings across all months for Austin, Boston, Miami, and San Francisco from 2019 to 2022.
city year corridor mean median sd
Austin 2019 destination 1.036 1.040 1.025
Austin 2020 destination 1.034 1.000 1.067
Austin 2021 destination 1.114 1.204 1.015
Austin 2022 destination 1.122 1.080 1.141
Austin 2019 origin 1.022 1.028 1.020
Austin 2020 origin 0.982 0.937 1.046
Austin 2021 origin 1.076 1.172 1.002
Austin 2022 origin 1.180 1.215 1.143
Boston 2019 destination 1.043 1.038 1.038
Boston 2020 destination 1.250 1.208 1.223
Boston 2021 destination 1.200 1.393 1.036
Boston 2022 destination 1.374 1.576 1.184
Boston 2019 origin 1.046 1.057 1.038
Boston 2020 origin 1.068 1.033 1.088
Boston 2021 origin 1.137 1.268 1.009
Boston 2022 origin 1.287 1.374 1.188
Miami 2019 destination 1.039 1.043 1.022
Miami 2020 destination 1.011 0.942 1.061
Miami 2021 destination 1.010 1.097 0.941
Miami 2022 destination 1.120 1.085 1.121
Miami 2019 origin 1.006 0.972 1.049
Miami 2020 origin 0.934 0.888 1.025
Miami 2021 origin 1.092 1.197 1.045
Miami 2022 origin 1.128 1.089 1.170
San Francisco 2019 destination 1.000 0.969 1.032
San Francisco 2020 destination 1.023 0.875 1.167
San Francisco 2021 destination 0.939 0.966 0.932
San Francisco 2022 destination 1.169 1.242 1.104
San Francisco 2019 origin 1.052 1.048 1.053
San Francisco 2020 origin 0.998 1.027 1.007
San Francisco 2021 origin 1.173 1.345 0.995
San Francisco 2022 origin 1.295 1.377 1.188
Table S5: Year-over-Year (YoY) metrics for domestic travel using 2018 as the base year. The table presents the mean, median, and standard deviation (sd) of lead times for Airbnb bookings across all months for Austin, Boston, Miami, and San Francisco from 2019 to 2022.
city year corridor mean median sd
Austin 2019 destination 1.008 1.013 1.004
Austin 2020 destination 1.281 1.261 1.230
Austin 2021 destination 0.791 0.748 0.838
Austin 2022 destination 0.829 0.776 0.887
Austin 2019 origin 1.029 1.009 1.052
Austin 2020 origin 1.263 1.125 1.304
Austin 2021 origin 0.880 0.851 0.942
Austin 2022 origin 1.060 1.068 1.064
Boston 2019 destination 1.057 1.049 1.065
Boston 2020 destination 1.567 1.741 1.381
Boston 2021 destination 1.060 1.090 1.033
Boston 2022 destination 1.159 1.257 1.072
Boston 2019 origin 1.019 1.014 1.022
Boston 2020 origin 1.256 1.153 1.270
Boston 2021 origin 0.876 0.827 0.962
Boston 2022 origin 1.063 1.064 1.070
Miami 2019 destination 1.006 0.991 1.025
Miami 2020 destination 1.288 1.270 1.269
Miami 2021 destination 0.678 0.605 0.845
Miami 2022 destination 0.827 0.744 0.940
Miami 2019 origin 0.999 0.985 1.010
Miami 2020 origin 1.081 0.944 1.170
Miami 2021 origin 0.865 0.864 0.921
Miami 2022 origin 1.020 1.019 1.029
San Francisco 2019 destination 1.020 1.013 1.040
San Francisco 2020 destination 1.389 1.441 1.313
San Francisco 2021 destination 0.811 0.751 0.946
San Francisco 2022 destination 0.945 0.975 0.941
San Francisco 2019 origin 1.041 1.029 1.050
San Francisco 2020 origin 1.169 1.034 1.232
San Francisco 2021 origin 0.877 0.865 0.933
San Francisco 2022 origin 1.101 1.130 1.081
Table S6: Year-over-Year (YoY) metrics for international travel using 2018 as the base year. The table presents the mean, median, and standard deviation (sd) of lead times for Airbnb bookings across all months for Austin, Boston, Miami, and San Francisco from 2019 to 2022.
city year corridor mean median q025 q0975
Austin 2019 destination 0.04 0.04 0.03 0.07
Austin 2020 destination 0.09 0.08 0.03 0.16
Austin 2021 destination 0.13 0.12 0.05 0.27
Austin 2022 destination 0.07 0.06 0.04 0.11
Austin 2019 origin 0.04 0.04 0.03 0.05
Austin 2020 origin 0.09 0.09 0.04 0.17
Austin 2021 origin 0.13 0.13 0.04 0.27
Austin 2022 origin 0.07 0.07 0.04 0.10
Boston 2019 destination 0.04 0.03 0.03 0.07
Boston 2020 destination 0.11 0.09 0.05 0.23
Boston 2021 destination 0.15 0.14 0.05 0.30
Boston 2022 destination 0.08 0.07 0.04 0.15
Boston 2019 origin 0.04 0.04 0.03 0.05
Boston 2020 origin 0.11 0.10 0.04 0.20
Boston 2021 origin 0.16 0.14 0.06 0.31
Boston 2022 origin 0.07 0.07 0.04 0.11
Miami 2019 destination 0.04 0.04 0.03 0.06
Miami 2020 destination 0.10 0.09 0.04 0.19
Miami 2021 destination 0.15 0.12 0.06 0.30
Miami 2022 destination 0.08 0.09 0.04 0.14
Miami 2019 origin 0.04 0.04 0.03 0.05
Miami 2020 origin 0.09 0.08 0.04 0.14
Miami 2021 origin 0.13 0.10 0.06 0.24
Miami 2022 origin 0.07 0.07 0.04 0.10
San Francisco 2019 destination 0.05 0.04 0.03 0.07
San Francisco 2020 destination 0.14 0.12 0.04 0.24
San Francisco 2021 destination 0.19 0.19 0.04 0.34
San Francisco 2022 destination 0.12 0.12 0.06 0.22
San Francisco 2019 origin 0.03 0.03 0.03 0.05
San Francisco 2020 origin 0.10 0.10 0.03 0.17
San Francisco 2021 origin 0.16 0.13 0.06 0.30
San Francisco 2022 origin 0.07 0.06 0.04 0.11
Table S7: $l^{1}_{{\text{YoY}},t}$: Year-over-Year normalized L1 divergence for domestic travel lead times, comparing the distributions of bookings from 2019 to 2022 for Austin, Boston, Miami, and San Francisco.
city year corridor mean median q025 q0975
Austin 2019 destination 0.10 0.09 0.07 0.14
Austin 2020 destination 0.25 0.26 0.09 0.41
Austin 2021 destination 0.34 0.32 0.15 0.60
Austin 2022 destination 0.14 0.13 0.09 0.25
Austin 2019 origin 0.05 0.05 0.05 0.06
Austin 2020 origin 0.21 0.17 0.05 0.47
Austin 2021 origin 0.24 0.17 0.07 0.59
Austin 2022 origin 0.10 0.10 0.07 0.13
Boston 2019 destination 0.05 0.05 0.04 0.07
Boston 2020 destination 0.31 0.33 0.06 0.59
Boston 2021 destination 0.37 0.39 0.11 0.67
Boston 2022 destination 0.13 0.13 0.06 0.19
Boston 2019 origin 0.04 0.04 0.03 0.05
Boston 2020 origin 0.21 0.19 0.04 0.45
Boston 2021 origin 0.25 0.17 0.07 0.57
Boston 2022 origin 0.11 0.10 0.07 0.15
Miami 2019 destination 0.04 0.04 0.04 0.05
Miami 2020 destination 0.28 0.28 0.04 0.47
Miami 2021 destination 0.37 0.27 0.14 0.69
Miami 2022 destination 0.13 0.13 0.07 0.20
Miami 2019 origin 0.04 0.04 0.03 0.05
Miami 2020 origin 0.14 0.13 0.04 0.30
Miami 2021 origin 0.16 0.11 0.06 0.38
Miami 2022 origin 0.08 0.08 0.06 0.11
San Francisco 2019 destination 0.06 0.06 0.05 0.07
San Francisco 2020 destination 0.32 0.34 0.06 0.55
San Francisco 2021 destination 0.40 0.42 0.11 0.70
San Francisco 2022 destination 0.15 0.13 0.09 0.24
San Francisco 2019 origin 0.04 0.03 0.03 0.05
San Francisco 2020 origin 0.18 0.16 0.04 0.38
San Francisco 2021 origin 0.21 0.13 0.07 0.49
San Francisco 2022 origin 0.11 0.11 0.05 0.18
Table S8: $l^{1}_{{\text{YoY}},t}$: Year-over-Year normalized L1 divergence for international travel lead times, comparing the distributions of bookings from 2019 to 2022 for Austin, Boston, Miami, and San Francisco.
city year corridor mean median q025 q0975
Austin 2019 destination 0.04 0.04 0.03 0.07
Austin 2020 destination 0.09 0.07 0.04 0.16
Austin 2021 destination 0.09 0.08 0.05 0.13
Austin 2022 destination 0.06 0.06 0.04 0.08
Austin 2019 origin 0.04 0.04 0.03 0.05
Austin 2020 origin 0.10 0.08 0.04 0.18
Austin 2021 origin 0.08 0.08 0.04 0.11
Austin 2022 origin 0.07 0.07 0.05 0.09
Boston 2019 destination 0.04 0.03 0.03 0.07
Boston 2020 destination 0.13 0.10 0.08 0.24
Boston 2021 destination 0.13 0.13 0.09 0.16
Boston 2022 destination 0.15 0.15 0.11 0.18
Boston 2019 origin 0.04 0.04 0.03 0.05
Boston 2020 origin 0.12 0.10 0.06 0.21
Boston 2021 origin 0.10 0.10 0.08 0.12
Boston 2022 origin 0.11 0.11 0.09 0.13
Miami 2019 destination 0.04 0.04 0.03 0.06
Miami 2020 destination 0.11 0.08 0.05 0.23
Miami 2021 destination 0.08 0.07 0.05 0.14
Miami 2022 destination 0.05 0.05 0.03 0.08
Miami 2019 origin 0.04 0.04 0.03 0.05
Miami 2020 origin 0.09 0.08 0.05 0.15
Miami 2021 origin 0.07 0.07 0.04 0.12
Miami 2022 origin 0.05 0.05 0.04 0.07
San Francisco 2019 destination 0.05 0.04 0.03 0.07
San Francisco 2020 destination 0.16 0.17 0.06 0.26
San Francisco 2021 destination 0.09 0.09 0.06 0.14
San Francisco 2022 destination 0.09 0.09 0.06 0.12
San Francisco 2019 origin 0.03 0.03 0.03 0.05
San Francisco 2020 origin 0.11 0.11 0.04 0.17
San Francisco 2021 origin 0.12 0.11 0.09 0.16
San Francisco 2022 origin 0.11 0.12 0.09 0.14
Table S9: $l^{1}_{{\text{YoY2018}},t}$: Normalized L1 divergence for domestic travel lead times using 2018 as the baseline year, comparing the distributions of bookings from 2019 to 2022 for Austin, Boston, Miami, and San Francisco.
city year corridor mean median q025 q0975
Austin 2019 destination 0.10 0.09 0.07 0.14
Austin 2020 destination 0.25 0.24 0.09 0.43
Austin 2021 destination 0.19 0.14 0.11 0.36
Austin 2022 destination 0.12 0.11 0.08 0.18
Austin 2019 origin 0.05 0.05 0.05 0.06
Austin 2020 origin 0.21 0.17 0.06 0.47
Austin 2021 origin 0.09 0.08 0.05 0.15
Austin 2022 origin 0.06 0.07 0.05 0.08
Boston 2019 destination 0.05 0.05 0.04 0.07
Boston 2020 destination 0.32 0.34 0.08 0.60
Boston 2021 destination 0.13 0.12 0.09 0.16
Boston 2022 destination 0.11 0.11 0.07 0.14
Boston 2019 origin 0.04 0.04 0.03 0.05
Boston 2020 origin 0.21 0.19 0.05 0.46
Boston 2021 origin 0.09 0.08 0.05 0.14
Boston 2022 origin 0.06 0.06 0.04 0.09
Miami 2019 destination 0.04 0.04 0.04 0.05
Miami 2020 destination 0.28 0.29 0.05 0.47
Miami 2021 destination 0.21 0.22 0.14 0.30
Miami 2022 destination 0.11 0.10 0.07 0.15
[…]